A trolley with mass m1 of 3.3 kg is connected to a light string with a weight of mass m2 of 2.2 kg.
The cord runs through a pulley (a small wheel where the cord runs freely)
a) Calculate the vehicle acceleration
b) Calculate the clamping force in the card
1
Expert's answer
2020-06-29T14:13:13-0400

As per the question,

mass of the trolley "(m_1)=3.3kg"

mass connected with the trolley with the help of the string "(m_2)=2.2kg"

chord runs through the pulley,

Let the acceleration of the trolley = a and tension in the string is T,

Now,

a)

"m_1g-T=m_1a"

"\Rightarrow 3.3g-T=3.3a----(i)"

and

"T-m_2g=m_2a"

"\Rightarrow T-2.2g=2.2a---(ii)"

from equation (i) and (ii),

"\Rightarrow 3.3g-2.2g=5.5a"

"\Rightarrow a=\frac{1.1g}{5.5}m/sec^2"

and

b)

"T=2.2g+2.2a=2.2g+2.2\times \frac{1.1g}{5.5}"

"\Rightarrow T=2.2g+0.44g=2.64gN"

Hence , clamping force on the card"=2T"

"=2\times 2.64g = 5.28gN"